1. The distribution of blood types in the United States is as follows:
O-positive: 38 %
O-negative: 7 %
A-positive: 34 %
A-negative: 6 %
B-positive: 9 %
B-negative: 2 %
AB-positive: 3 %
AB-negative: 1 %
a. What percentage of the population is Rh positive?

b. Given that a person is Rh negative, what is the probability that he or she will be type 0, a universal donor?


c. If you randomly select 15 people from the U.S. population, what is the probability that at least one person will have type AB blood?


d. Would you expect a similar result if you sampled 15 people from Asia? Explain

Answer by mathmate(429)

Question:
1. The distribution of blood types in the United States is as follows:
O-positive: 38 %
O-negative: 7 %
A-positive: 34 %
A-negative: 6 %
B-positive: 9 %
B-negative: 2 %
AB-positive: 3 %
AB-negative: 1 %
a. What percentage of the population is Rh positive?
b. Given that a person is Rh negative, what is the probability that he or she will be type 0, a universal donor?
c. If you randomly select 15 people from the U.S. population, what is the probability that at least one person will have type AB blood?
d. Would you expect a similar result if you sampled 15 people from Asia? Explain

Solution:
(a) Percentage of Rh+ =(38+34+9+3)=84%

(b)
P(O|+)
=P(O∩+)/P(+)
=0.38/0.84
=19/42
(=p.4524, approx.)

(c)
We calculate the probability of no AB from a sample of 15, and subtract from 1.0 to get P(at least one AB).
P(no AB)=1-(0.03+0.01)=0.96
P(15 no AB)=0.96^15=0.5421 (approx.)
=>
P(at least one AB out of 15)
=1-0.5421
=0.4579

(d) no, won't expect the same proportions. Percentages vary greatly with geography.